# Imitation Learning from a Single Temporally Misaligned Video

### v1

@@ 6 Related Works
Learning From a Few Video Demonstrations. Some works focus on training a reward model that takes videos as input and outputs scalar rewards for RL (Sontakke et al., 2024; Yang et al., 2024), but these models often require fine-tuning on task-specific data to improve performance (Fu et al., 2024a). Instead, we follow IRL’s formulation where we aim to match the learner and demonstration distributions, which is equivalent to optimizing the Integral Probabitliy Metrics (IPMs) (Sun et al., 2019; Swamy et al., 2021).
Optimal Transport Used In IRL. Recent works leverage OT (Peyré & Cuturi, 2020) to optimize IPMs. In the imitation learning setting, where a teleoperated demonstration dataset contains state-action labels, prior works can directly minimize the Wasserstein distance between the learner and the demonstration’s state-action distributions (Xiao et al., 2019; Dadashi et al., 2020; Papagiannis & Li, 2022; Luo et al., 2023; Bobrin et al., 2024). These approaches do not work given only a single visual demonstration. Without access to privileged states, recent works instead uses a distance function that measures the transport cost between two visual embeddings (Cohen et al., 2022; Haldar et al., 2023b, a; Guzey et al., 2024; Tian et al., 2024; Liu et al., 2024; Fu et al., 2024b; Kedia et al., 2024). Both types of approaches assume that a Markovian function measuring transport cost between two timesteps is sufficient. However, our work tackles sequence-matching problems, where agents must follow the subgoals from a temporally misaligned demonstration in the correct order. OT with a Markovian function fails because the true reward function now depends on the entire trajectory, a limitation that our approach addresses.
@@ 7 Discussion
We investigate sequence-matching tasks, where the learner must follow a expert video demonstration that may be temporally misaligned. We analyze how algorithms that match the learner and expert distribution at the frame level (e.g., optimal transport) result in reward functions that fail in this setting. Following our key insight that matching should be defined at the sequence level, we present ORCA: a principled reward function that computes the probability that the learner has covered every subgoal in the correct order. Experiments on Meta-world and Humanoid tasks show that ORCA rewards train agents that complete the tasks efficiently regardless of the level of temporal misalignment. We recognize a few limitations: (1) ORCA relies on a good visual distance metric that measures the similarity between two frame embeddings. Future work will explore using online finetuning to improve the encoder (Fu et al., 2024a) and solve cross-embodiment tasks where temporal misalignment is common. (2) If the demonstration contains unrealizable subgoals that the policy can never achieve (e.g., due to physical limitations), the policy will fail. We plan to explore how ORCA can help iteratively identify realizable subgoals.
@@ A.1 Proofs of ORCA Desiderata
Let $\xi^{-}$ be a trajectory that is out of order; specifically, there exists a subgoal $\tilde{o}_{j}$ such that $\tilde{o}_{j}$ is occupied at time $t$ and $\tilde{o}_{j-1}$ is not yet covered. Let $\xi^{+}$ be a trajectory that is identical to $\xi^{-}$, except that it covers $\tilde{o}_{j-1}$ before time $t$. Then, $\mathcal{R}_{ORCA}(o^{+}_{t},\tilde{\xi})>\mathcal{R}_{ORCA}(o^{-}_{t},\tilde{\xi})$.
@@
$\mathcal{R}_{ORCA}(o^{+}_{t},\tilde{\xi})>\mathcal{R}_{ORCA}(o^{-}_{t},\tilde{\xi}).$ (22)
@@
Let $\xi^{-}$ be a trajectory that occupies $\tilde{o}_{j-1}$ at time $t-1$ and continues to occupy $\tilde{o}_{j-1}$ at time $t$, instead of progressing towards $\tilde{o}_{j}$. Let $\xi^{+}$ be an identical trajectory that progresses towards $\tilde{o}_{j}$ at $t$, and assume that neither trajectory has been closer to $\tilde{o}_{j}$ before. Then, $\mathcal{R}_{ORCA}(o^{+}_{t},\tilde{\xi})>\mathcal{R}_{ORCA}(o^{-}_{t},\tilde{\xi})$.
@@
$\mathcal{R}_{ORCA}(o^{+}t,\tilde{\xi})>\mathcal{R}_{ORCA}(o^{-}_{t},\tilde{\xi}).$ (29)
@@ Appendix B Environment Details
We run experiments in two different environments: Meta-world (Yu et al., 2021), a manipulation environment, and Humanoid-v4 (Todorov et al., 2012), a more difficult control environment. Following prior work (Fu et al., 2024b), we test 10 different tasks in Meta-world. To show that ORCA works in more general domains, we additionally design 4 tasks in the humanoid environment that require moving the arms of the humanoid.
@@ B.1 Visual Encoder
In Meta-world, we follow prior work (Fu et al., 2024b) in using an off-the-shelf Resnet50 (He et al., 2015) as the visual encoder. We use the cosine distance between embeddings to produce a cost matrix.
@@ B.2.1 Tasks
We run experiments on 10 different tasks in the Meta-world (Yu et al., 2021) environment. In addition to the 9 tasks in Fu et al. (2024b), we added Door-close. We classify them into easy, medium, and hard based on factors like their visual difficulty, necessity for precise motor control, and interaction with other objects. For further information on the tasks, we refer the reader to (Fu et al., 2024b). Below is a brief description of the objective for each task:
@@ B.2.2 Training Details
1) Context embedding: We use the context embedding-based cost matrix proposed in (Fu et al., 2024b), which can be interpreted as a diagonal smoothing kernel. Specifically, the distance between two frames is expressed as the average distance over the next $c_{w}$ learner and demonstration frames (where $c_{w}$ refers to the context window):
@@
We choose a context window of length 3, which resulted in the best performance in (Fu et al., 2024b). Although a longer context window could damage performance on extremely mismatched tasks, we find that a small window helps regularize the noisiness of the visual distance metric.
@@ B.3.1 Tasks
The humanoid’s goal is to follow the motion of a demonstration trajectory within a maximum of 120 timesteps. We define 4 motions, corresponding to 4 demonstration trajectories. The humanoid must remain standing while doing all tasks. The visual demonstrations for each task are shown in 9.
@@
This distance metric is used for by the distribution-matching algorithms and ORCA. Then, we compute the confidence of the learner observation embedding $c(\phi(o^{L}_{t}))$ according to 33. This results in the final reward function:
$\mathcal{R}=c(\phi(o^{L}_{t}))*\mathcal{R}_{seq}(o_{t}^{L},\xi^{D})+\lambda\mathcal{R}_{stability}$ (32)
@@ B.3.3 Joint Predictor Training Details
Training: To train our joint predictor $f\circ\phi$, we fully fine-tune a ResNet50 backbone (He et al., 2015) pre-trained on ImageNet-1K (Deng et al., 2009) with a 3-layer MLP head that projects to the joint dimension. The MLP head has layers of shape (2048, 1024), (1024, 1024), and (1024, 54), where 54 represents the number of joints (18) multiplied by the dimension per joint (3). Optimization is performed over 100 epochs using SGD with learning rate .008, batch size 16, and momentum 0.875. After training the joint predictor, we freeze the backbone weights, and train a shallow autoencoder architecture with two linear layers of shapes $(d_{\phi},32)$ and $(32,d_{\phi})$ using the same parameters, where $d_{\phi}$ is the dimension of the backbone (2048 in this case). This provides the reconstruction loss that is used for confidence estimation, as described in B.3.4
@@ B.3.4 Confidence Estimate as Reward Scaling
Appendix C Experiments
C.1 RL Policy Details.
In the Meta-world environment, we use an identical setup and hyperparameters to Fu et al. (2024b), training the policy with DrQ-v2 over 1 million steps. We also follow Fu et al. (2024b) to repeat a policy’s predicted action 2 times, which effectively shortens the episode length by half. This effect is applied both in RL training and generating demonstrations from hand-engineered policies. In MuJoCo, the policy is trained with SAC over 2 million steps, and we slightly modify the parameters from Rocamonde et al. (2024) to adapt to the shorter training period (2 million steps vs 10 million steps). All policies are trained with ground truth state observations, but we emphasize that the reward function only sees visual observations.
@@ Appendix C Experiments
  • TemporalOT. TemporalOT (Fu et al., 2024b) was originally designed for learner and reference sequences of the same length. In this paper, we compare against a slightly stronger version of TemporalOT, which allows for a learner that is linear in the speed of the expert (either faster or slower). Specifically, we mask with a windowed diagonal matrix stretched along the longer of the learner and demonstration axis. Because TemporalOT also formulates the entropic regularized OT problem (4), the entropic weight is also $\epsilon=1$. We use a varying mask window size, depending on the reference length. Fu et al. (2024b) used $k_{m}=10$ for metaworld matched demos. We use the same for matched demos, and for mismatched demos, we let $k_{m}\approx\lceil\frac{|\tilde{\xi}|}{10}\rceil$.
@@
C.3 Varying Temporal Alignment Level Experiment Details.
We study the effect of how increasingly misaligned demonstrations would affect ORCA and TemporalOT’s performances. We identify two types of temporal misalignment: either the demonstration contains pauses and is slower, or the demonstration accelerates through a segment and is faster. We randomly perturb the original demonstrations of three Meta-world tasks (Door-open, Window-open, Lever-pull), generating 6 demonstrations per task per speed type (i.e, slow or fast). Concretely, we first evenly split the original demonstration into five segments. To get the first three randomly perturbed demonstrations, we randomly select one segment and randomly change their speed. We can speed up a segment by 2, 4, 6, 8, or 10 times, and we can slow down a segment by 2, 3, 4, 5, or 6 times. To get the rest of the randomly perturbed demonstrations, we randomly select three segments and randomly change their speed. Then, for each task and speed type, we categorize 3 of the perturbed demonstrations as having “Low” level of misalignment and the other 3 as having “High” level of misalignment by ranking the demonstrations based on the mean absolute deviation between the segment length. The more varied the segment lengths are with respect to each other, the more temporally misaligned this demonstration is. Each perturbed demonstration is trained on a random seed, and we ensure that ORCA and TemporalOT are trained on that same seed for fair comparison.
C.4 Additional Experiment Results
@@
Fig. 14 shows a qualitative comparison between frame-level matching approaches and ORCA on the left arm up task. ORCA quickly solves the task, while the other approaches have the failure modes described in 3.
@@
C.5 Frame-Level Failure Modes in Meta-world
@@
OT and TemporalOT Fail to Enforce Ordering.
Fig. 16 shows how OT and TemporalOT with a large $k_{m}$ both fail to enforce subgoal ordering, thus rewarding trajectories that do not complete subgoals in the correct order.
@@
DTW Fails to Enforce Subgoal Coverage.
Fig. 17 shows how DTW fails to enforce full subgoal coverage, getting stuck in an intermediate subgoal.
@@
C.6 The importance of pretraining for ORCA.
@@
Figure 18: Example of pretraining improve ORCA’s performance.

### v2

@@ 6 Related Works
Learning From a Few Video Demonstrations. Prior works in imitation learning (Jain et al., 2024; Fu et al., 2024a; Xu et al., 2022; Duan et al., 2017; Palo & Johns, 2024) train a policy that, at inference time, takes a video-action demonstration and the current robot state to output actions. This is different from our policy formulation because ORCA and its baselines do not have access to demonstrations with action labels. Closer to our setting, some works focus on training a reward model that takes videos as input and outputs scalar rewards for RL (Sontakke et al., 2024; Yang et al., 2024), but these models often require fine-tuning on task-specific data to improve performance (Fu et al., 2024b). We follow IRL’s formulation where we aim to match the learner and demonstration distributions, which is equivalent to optimizing the Integral Probability Metrics (IPMs) (Sun et al., 2019; Swamy et al., 2021).
Optimal Transport Used In IRL. Recent works leverage OT (Peyré & Cuturi, 2020) to optimize IPMs. In the imitation learning setting, where a teleoperated demonstration dataset contains state-action labels, prior works can directly minimize the Wasserstein distance between the learner and the demonstration’s state-action distributions (Xiao et al., 2019; Dadashi et al., 2021; Papagiannis & Li, 2022; Luo et al., 2023; Bobrin et al., 2024). These approaches do not work given only a single visual demonstration. Without access to privileged states, recent works instead use a distance function that measures the transport cost between two visual embeddings (Cohen et al., 2022; Haldar et al., 2023b, a; Guzey et al., 2024; Tian et al., 2024; Liu et al., 2024; Fu et al., 2024c; Kedia et al., 2025). Both types of approaches assume that a Markovian function measuring transport cost between two timesteps is sufficient. However, our work tackles sequence-matching problems, where agents must follow the subgoals from a temporally misaligned demonstration in the correct order. OT with a Markovian distance function fails because the true reward function depends on the entire trajectory, a limitation that our approach addresses.
@@ 7 Discussion
We investigate sequence-matching tasks, where the learner must follow a expert video demonstration that may be temporally misaligned. We analyze how algorithms that match the learner and expert distribution at the frame level (e.g., optimal transport) result in reward functions that fail in this setting. Following our key insight that matching should be defined at the sequence level, we present ORCA: a principled reward function that computes the probability that the learner has covered every subgoal in the correct order. Experiments on Meta-world and Humanoid tasks show that ORCA rewards train agents that complete the tasks efficiently regardless of the level of temporal misalignment.
We recognize a few limitations of ORCA: (1) it relies on a good visual distance metric that measures the similarity between two frames. Future work will explore using online finetuning to improve the encoder (Fu et al., 2024b) and solve cross-embodiment tasks where temporal misalignment is common. (2) video-following may result in unexpected failures due to task misspecification. We plan to explore how ORCA can be applied to language subgoals in addition to video frames.
Acknowledgments
This work was supported in part by the National Science Foundation FRR (#2327973).
@@ A.1 Proofs of ORCA Desiderata
Let $\xi^{-}$ be a trajectory that is out of order; specifically, there exists a subgoal $\tilde{o}_{j}$ such that $\tilde{o}_{j}$ is occupied at time $t$ and $\tilde{o}_{j-1}$ is not yet covered. Let $\xi^{+}$ be a trajectory that is identical to $\xi^{-}$, except that it covers $\tilde{o}_{j-1}$ before time $t$. Then, $\mathcal{R}_{ORCA}(o^{+}_{t},\xi^{+},\tilde{\xi})>\mathcal{R}_{ORCA}(o^{-}_{t},\xi^{-},\tilde{\xi})$.
@@
$\mathcal{R}_{ORCA}(o^{+}_{t},\xi^{+},\tilde{\xi})>\mathcal{R}_{ORCA}(o^{-}_{t},\xi^{-},\tilde{\xi}).$ (23)
@@
Let $\xi^{-}$ be a trajectory that occupies $\tilde{o}_{j-1}$ at time $t-1$ and continues to occupy $\tilde{o}_{j-1}$ at time $t$, instead of progressing towards $\tilde{o}_{j}$. Let $\xi^{+}$ be an identical trajectory that progresses towards $\tilde{o}_{j}$ at $t$, and assume that neither trajectory has been closer to $\tilde{o}_{j}$ before. Then, $\mathcal{R}_{ORCA}(o^{+}_{t},\xi^{+},\tilde{\xi})>\mathcal{R}_{ORCA}(o^{-}_{t},\xi^{-},\tilde{\xi})$.
@@
$\mathcal{R}_{ORCA}(o^{+}_{t},\xi^{+},\tilde{\xi})>\mathcal{R}_{ORCA}(o^{-}_{t},\xi^{-},\tilde{\xi}).$ (30)
@@ Appendix B Environment Details
We run experiments in two different environments: Meta-world (Yu et al., 2020), a manipulation environment, and Humanoid-v4 (Todorov et al., 2012), a more difficult control environment. Following prior work (Fu et al., 2024c), we test 10 different tasks in Meta-world. To show that ORCA works in more general domains, we additionally design 4 tasks in the humanoid environment that require moving the arms of the humanoid.
@@ B.1 Visual Encoder
In Meta-world, we follow prior work (Fu et al., 2024c) in using an off-the-shelf Resnet50 (He et al., 2016) as the visual encoder. We use the cosine distance between embeddings to produce a cost matrix.
@@ B.2.1 Tasks
We run experiments on 10 different tasks in the Meta-world (Yu et al., 2020) environment. In addition to the 9 tasks in Fu et al. (2024c), we added Door-close. We classify them into easy, medium, and hard based on factors like their visual difficulty, necessity for precise motor control, and interaction with other objects. For further information on the tasks, we refer the reader to (Fu et al., 2024c). Below is a brief description of the objective for each task:
@@ B.2.2 Training Details
1) Context embedding: We use the context embedding-based cost matrix proposed in (Fu et al., 2024c), which can be interpreted as a diagonal smoothing kernel. Specifically, the distance between two frames is expressed as the average distance over the next $c_{w}$ learner and demonstration frames (where $c_{w}$ refers to the context window):
@@
We choose a context window of length 3, which resulted in the best performance in (Fu et al., 2024c). Although a longer context window could damage performance on extremely mismatched tasks, we find that a small window helps regularize the noisiness of the visual distance metric.
@@ B.3.1 Tasks
The humanoid’s goal is to follow the motion of a demonstration trajectory within a maximum of 120 timesteps. We define 4 motions, corresponding to 4 demonstration trajectories. The humanoid must remain standing while doing all tasks. The visual demonstrations for each task are shown in 10.
@@
This distance metric is used by ORCA to obtain $\mathcal{R}_{orca}$. Then, we compute the confidence of the learner observation embedding $c(\phi(o^{L}_{t}))$ according to 34. This results in the final reward function:
$\mathcal{R}=c(\phi(o^{L}_{t}))*\mathcal{R}_{orca}(o_{t}^{L},\xi^{D})+\lambda\mathcal{R}_{stability}$ (33)
@@ B.3.3 Joint Predictor Training Details
Training: To train our joint predictor $f\circ\phi$, we fully fine-tune a ResNet50 backbone (He et al., 2016) pre-trained on ImageNet-1K (Deng et al., 2009) with a 3-layer MLP head that projects to the joint dimension. The MLP head has layers of shape (2048, 1024), (1024, 1024), and (1024, 54), where 54 represents the number of joints (18) multiplied by the dimension per joint (3). Optimization is performed over 100 epochs using SGD with learning rate .008, batch size 16, and momentum 0.875. After training the joint predictor, we freeze the backbone weights, and train a shallow autoencoder architecture with two linear layers of shapes $(d_{\phi},32)$ and $(32,d_{\phi})$ using the same parameters, where $d_{\phi}$ is the dimension of the backbone (2048 in this case). This provides the reconstruction loss that is used for confidence estimation, as described in B.3.4
@@ B.3.4 Confidence Estimate as Reward Scaling
B.4 RL Policy.
In the Meta-world environment, we use an identical setup and hyperparameters to Fu et al. (2024c), training the policy with DrQ-v2 over 1 million steps. We also follow Fu et al. (2024c) to repeat a policy’s predicted action 2 times, which effectively shortens the episode length by half. This effect is applied both in RL training and generating demonstrations from hand-engineered policies. In MuJoCo, the policy is trained with SAC over 2 million steps, and we slightly modify the parameters from Rocamonde et al. (2024) to adapt to the shorter training period (2 million steps vs 10 million steps). All policies are trained with ground truth state observations, but we emphasize that the reward function only sees visual observations.
@@
  • TemporalOT. TemporalOT (Fu et al., 2024c) was originally designed for learner and reference sequences of the same length. In this paper, we compare against a slightly stronger version of TemporalOT, which allows for a learner that is linear in the speed of the expert (either faster or slower). Specifically, we mask with a windowed diagonal matrix stretched along the longer of the learner and demonstration axis. Because TemporalOT also formulates the entropic regularized OT problem (4), the entropic weight is also $\epsilon=1$. We use a varying mask window size, depending on the reference length. Fu et al. (2024c) used $k_{m}=10$ for metaworld matched demos. We use the same for matched demos, and for mismatched demos, we let $k_{m}\approx\lceil\frac{|\tilde{\xi}|}{10}\rceil$.
@@
B.6 Varying Temporal Alignment Level Experiment Details.
We study the effect of how increasingly misaligned demonstrations would affect ORCA and TemporalOT’s performances. We identify two types of temporal misalignment: either the demonstration contains pauses and is slower, or the demonstration accelerates through a segment and is faster. We randomly perturb the original demonstrations of three Meta-world tasks (Door-open, Window-open, Lever-pull), generating 6 demonstrations per task per speed type (i.e, slow or fast).
Concretely, we first evenly split the original demonstration into five segments. To get the first three randomly perturbed demonstrations, we randomly select one segment and randomly change their speed. We can speed up a segment by 2, 4, 6, 8, or 10 times, and we can slow down a segment by 2, 3, 4, 5, or 6 times. To get the rest of the randomly perturbed demonstrations, we randomly select three segments and randomly change their speed. Then, for each task and speed type, we categorize 3 of the perturbed demonstrations as having “Low" level of misalignment and the other 3 as having “High" level of misalignment by ranking the demonstrations based on the mean absolute deviation between the segment length. The more varied the segment lengths are with respect to each other, the more temporally misaligned this demonstration is. Each perturbed demonstration is trained on a random seed, and we ensure that ORCA and TemporalOT are trained on that same seed for fair comparison.
Appendix C Additional Experiment Results
@@
Fig. 15 shows a qualitative comparison between frame-level matching approaches and ORCA on the left arm up task. ORCA quickly solves the task, while the other approaches have the failure modes described in Sec. 3.
@@
C.1 Frame-Level Failure Modes in Meta-world
@@
OT and TemporalOT Fail to Enforce Ordering: Fig. 17 shows how OT and TemporalOT with a large $k_{m}$ both fail to enforce subgoal ordering, thus rewarding trajectories that do not complete subgoals in the correct order.
@@
DTW Fails to Enforce Subgoal Coverage: Fig. 18 shows how DTW fails to enforce full subgoal coverage, getting stuck in an intermediate subgoal.
@@
C.2 The importance of pretraining for ORCA
@@
Figure 19: Example of how pretraining improves ORCA’s performance.
C.3 Additional Baselines
In addition to the baselines described in the main paper, we include a classic IRL algorithm and a text-conditioned reward function. We test these baselines on all easy and medium Meta-world tasks. The results are shown in Table 6.
GAIfO (Torabi et al., 2019) is a traditional IRL algorithm. It jointly trains an RL policy and a discriminator between state transitions from the learner and demonstration distributions. The discriminator is used as the reward function for the policy, encouraging the agent to stay within the expert state distribution. Due to computation limitations, we used a state-based demonstration instead of a video. In contrast, ORCA and all existing baselines use video demonstrations. GAIfO outperforms TemporalOT, but it is still significantly worse than ORCA. We hypothesize that, because there is only one demonstration, GAIfO could fixate on inconsequential details instead of estimating the task reward.
LIV (Ma et al., 2023) is a vision language model that can be used to calculate the rewards of a video that teaches given a description of text. For this baseline, we follow the implementation of FuRL [3], which closely resembles our IRL setup. While it outperforms TemporalOT, LIV is significantly worse than ORCA. LIV uses a single text goal to describe the task, whereas ORCA is conditioned on a sequence of images, making ORCA denser and better at capturing details. Future work should explore how ORCA can be applied to sequences of text subgoals.
Table 6: Meta-world per-task results on additional baselines LIV and GAIfO. We report the mean expert-normalized return and the standard error across different methods and task difficulties.
Category Task GAIfO (state-based demo) LIV (text goal) TemporalOT ORCA
Easy Button-press 0.12 (0.04) 0.00 (0.00) 0.10 (0.02) 0.62 (0.11)
Easy Door-close 0.34 (0.04) 0.34 (0.09) 0.19 (0.01) 0.88 (0.01)
Medium Door-open 0.22 (0.05) 0.00 (0.00) 0.08 (0.01) 0.89 (0.13)
Medium Window-open 0.27 (0.09) 0.72 (0.19) 0.26 (0.05) 0.85 (0.16)
Medium Lever-pull 0.10 (0.03) 0.00 (0.00) 0.07 (0.03) 0.28 (0.09)
Total 0.21 (0.03) 0.21 (0.05) 0.14 (0.01) 0.70 (0.05)
C.4 Visual Encoder Ablation
The choice of encoder is a practical, task-dependent choice, and not the main focus of this work. In Metaworld, we followed Fu et al. (2024c) and used a pretrained Resnet50. In Humanoid, we finetuned the encoder on a set of images from the simulation environment.
We include in 7 an ablation of ORCA with LIV (Ma et al., 2023), a robotics-specific visual encoder, and DINOv2 (Oquab et al., 2023), a standard vision model. Overall, Resnet50 achieves the best performance, although there is high variability.
Table 7: Comparison of different visual backbones with ORCA. We report the mean expert-normalized return and the standard error across different backbone architectures with their parameter counts.
Task ORCA+Resnet50 (26M) ORCA+LIV (100M) ORCA+DINOv2-L (300M)
Door-open 1.71 (0.08) 1.10 (0.06) 0.44 (0.12)
Window-open 0.50 (0.14) 1.17 (0.15) 0.65 (0.10)
Lever-pull 0.28 (0.09) 0.04 (0.01) 0.16 (0.06)
Total 0.83 (0.09) 0.77 (0.08) 0.42 (0.06)
C.5 Image-Conditioned Policy
ORCA’s goal is to estimate rewards from a single demonstration in the observation space before learning a policy, and we pose no assumptions on the policy itself. In this paper, we choose to provide the policy with ground truth states, for better comparisons with prior work. To investigate an image-conditioned policy, we trained an agent using DrQv2 [10] on a subset of the Metaworld tasks. Table 8 shows the results. The state and image-based policies trained with ORCA have similar performance (average normalized return: 0.704 ± 0.10 → 0.76 ± 0.06). In contrast, policies trained with TemporalOT perform poorly, regardless of their input.
Table 8: Comparison of policies trained on ground truth states and image observations. We report the mean expert-normalized return and the standard error across different policy types.
Task TemporalOT (state) TemporalOT (image) ORCA (state) ORCA (image)
Button-press 0.10 (0.02) 0.00 (0.00) 0.62 (0.11) 0.60 (0.11)
Door-close 0.19 (0.01) 0.12 (0.01) 0.88 (0.01) 0.88 (0.02)
Door-open 0.08 (0.01) 0.00 (0.00) 0.89 (0.13) 1.31 (0.12)
Window-open 0.26 (0.05) 0.28 (0.05) 0.85 (0.16) 0.79 (0.15)
Lever-pull 0.07 (0.03) 0.00 (0.00) 0.28 (0.09) 0.19 (0.08)
Total 0.14 (0.02) 0.08 (0.01) 0.71 (0.10) 0.76 (0.06)
C.6 Runtime Comparison of Approaches
We experimentally tested the reward calculation latency to validate that ORCA is similar in runtime to its baselines. For each method, we evaluated 100 rollouts with a demonstration of length 100 and learner rollouts of length 100 and 300.
Table 9 shows the average latency (ms) of different reward computation methods. ORCA is 24% faster than TemporalOT and 52% faster than RoboCLIP given rollouts of length 100, and it is comparable to threshold, OT, and DTW. LIV is the fastest method because it only considers a text goal and not a sequence of subgoals, but this leads to poor performance on most tasks. The ordering of the methods is the same for 100 and 300 frames, demonstrating that ORCA can efficiently scale to longer horizon tasks.
Table 9: Latency comparison of ORCA and its baselines. Methods are sorted by their latency in increasing order from left to right. We report the mean performance and standard error given learner rollouts of length 100 and 300.
Learner Length LIV OT Threshold ORCA DTW TemporalOT Roboclip
100 frames 38.0 (4.7) 54.0 (0.4) 54.7 (3.8) 56.9 (0.3) 58.8 (0.4) 75.1 (1.2) 118.2 (17.9)
300 frames 130.0 (3.3) 170.2 (0.4) 171.0 (4.2) 179.9 (0.6) 185.3 (0.4) 228.5 (2.0) 300.2 (31.3)